OpenStudy (anonymous):

ABCD,AB=2x+5,CD=2y+1,AD=y+5 and BC=3x-4 find x and y

OpenStudy (anonymous):

please answer it :)

OpenStudy (mathstudent55):

What is ABCD, a quadrilaterl, a trapezoid, a parallelogram, a rhombus, a rectangle, or a square? Is there a symbol before the letters ABCD?

OpenStudy (anonymous):

ABCD is a parallelogram

OpenStudy (anonymous):

no thats the only the question

OpenStudy (mathstudent55):

Ok, now that we know it's a parallelogram, what do you know about the sides of a parallelogram?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Look at figure, AB and CD are opposite sides. AD and BC are opposite sides. In a parallelogram, opposite sides are congruent. That means the lengths of opposite sides are equal.

OpenStudy (anonymous):

how about the formula?

OpenStudy (mathstudent55):

Set two equations. Each one is one side equal to the opposite side. Since AB and CD have the same length, you get: 2x + 5 = 2y + 1 Since AD and BC have same length, you get: y + 5 = 3x - 4 Now solve the two equations as a system of equations.

OpenStudy (anonymous):

i dont know it..please solve the answer :)

OpenStudy (anonymous):

do you know it:)?

OpenStudy (mathstudent55):

Ok, a system of equations: 2x + 5 = 2y + 1 y + 5 = 3x - 4 On first equation, subtract from both sides: 2x = 2y - 4 Now divide both sides by 2: x = y - 2 Now taks x = y - 2, and substitute it for x in the second original equation: y + 5 = 3(y - 2) - 4 Distribute 3 on right side: y + 5 = 3y - 6 - 4 Do subtraction on right side: y + 5 = 3y - 10 Now subtract 5 from both sides, and subtract 3y from both sides: -2y = -15 y = 7.5

OpenStudy (mathstudent55):

Now that we know y = 7.5, substitute it into the first equation: 2x + 5 = 2(7.5) + 1 2x + 5 = 15 + 1 2x + 5 = 16 2x = 11 x = 5.5